says that when the line 4x + 3y-24 is dilated by a scale factor of 2 centered at the point (3-4), the equation of the dlated ine is16 Is Alyah The use of the set of axes below is optional.] correct? Explain why
Answer #1

Note that if center of dilation is on the line then dilated line and the normal line both are on the same line....

Here (3,4) is on line 4x+3y = 24 ....

So dilated line is on the same line 4x+3y = 24 ...

But the given dilated line is 4x+3y = 16 ..this is not the dilated line becuase this line is parallel to the original line...

So allyah is not correct....